# Exploring the Frontiers of Authorship Attribution in The Era of LLMs

### v1

@@ 5. Human-LLM Co-authored Text Attribution
Besides creating text from scratch, LLMs are often used for extending sequences from human prompts. These perturbations, have diminished the effectiveness of existing text detection methods (Bhat and Parthasarathy, 2020b). Identifying text that combines input from both human authors and LLMs presents unique challenges. Hybrid texts may originate as human-written content, with LLMs employed for conditional generation, making it difficult to clearly distinguish between the stylistic features of human and machine contributions. There are fewer studies on this task due to its difficulty, and existing research often makes simplifications.
@@ 5.2. Methodologies
To effectively analyze and classify human-LLM co-authored texts versus those solely authored by humans or LLMs, feature-based methods from LLM-generated text detection, such as Log-likelihood (Solaiman et al., 2019), GLTR (Gehrmann et al., 2019a), and log-rank (Mitchell et al., 2023), are adapted to this task. Additionally, neural network-based models like BERT (Ippolito et al., 2020), Radar (Hu et al., 2023), and GPT-sentinel (Chen et al., 2023a) can also be applied. The complexity of this task increases as users may employ multiple LLMs to compose different sections of an article, further blurring the lines between human and machine-generated content. Consequently, techniques used in earlier LLM detection need to evolve continuously. This ongoing evolution in detection strategies mirrors the increasing sophistication of LLM outputs and the collaborative nature of modern text creation.
@@ 5.3. Open Challenges
Human-LLM co-authored texts pose a more intricate challenge due to the blending of stylistic and linguistic elements from both human authors and LLMs. These texts may begin as human drafts later extended or revised by LLMs, or vice versa, resulting in an integration of styles that obscures distinct authorial markers (Liu et al., 2023b). Advanced techniques are required to dissect these integrations, identifying where and how LLM contributions intersect with human input (Wang et al., 2024). This involves detecting subtle shifts in style and contextual cues that indicate the extent of LLM involvement, allowing for accurate segmentation and attribution of authorship within hybrid documents.
@@ 7.2. Generalization
This subsection examines the applicability of current methodologies across varying LLMs, domains, genres, and languages. Domains refer to broad areas of knowledge or topics, while genres refer to specific styles or forms of writing within any domain. Domain generalization poses significant challenges due to variability in vocabularies, syntax, and styles across different subjects, complicating accurate authorship attribution. Attribution performance tends to drop when known and query texts differ in topic or genre (Altakrori et al., 2021). Models like BERT and RoBERTa have shown limitations in cross-domain tasks (Barlas and Stamatatos, 2020b; Huertas-Tato et al., 2022), and adapting models to new domains remains difficult due to factors like dataset size variability and writing styles. Traditional methods focused on identifying less topic-dependent features such as function words and part-of-speech n-grams (Madigan et al., 2005b; Menon and Choi, 2011), while recent approaches highlight the importance of training more powerful transformer-based models (Rivera-Soto et al., 2021) and techniques such as adversarial training (Li et al., 2017; Ben-David et al., 2010; Ganin et al., 2016).
@@
Improving generalization can potentially be achieved through several strategies. First, leveraging transfer learning by pre-training on large, diverse datasets and fine-tuning on specific domains enhances adaptability and performance (Barlas and Stamatatos, 2020a; Rodriguez et al., 2022b). Second, developing domain and genre-invariant features would improve robustness by focusing on core stylistic elements (Argamon et al., 2003). Third, employing hybrid models and ensemble methods that integrate domain-specific knowledge can optimize prediction accuracy by drawing on the strengths of individual models (Bacciu et al., 2019). Additionally, incorporating contextual factors like the writing environment or intended audience, alongside data augmentation techniques, can bolster generalization. Finally, improving attribution in multilingual contexts enables models to operate effectively across various languages (Chen et al., 2022a; Shamardina et al., 2022). As detectors could be biased against non-native English writers (Liang et al., 2023), enhancing multilingual generalization is crucial for fairness. Collectively, these approaches foster robust and adaptable models equipped to handle diverse styles and contexts.
@@ 7.3. Explainability
Improving explainability is crucial for ensuring transparency and trust in authorship attribution models, especially as they become increasingly integrated into fields such as forensic investigation, academia, and journalism. Developing explainable AI techniques specifically for authorship attribution can lead to methodologies where the reasoning behind attributions is clear and understandable. This transparency is particularly important when authorship attribution is used as evidence in legal proceedings (Chaski, 2005; Rocha et al., 2016). Traditional attribution methods, which rely on stylistic and linguistic features to identify an author, often struggle to distinguish between human-authored texts and those generated by large language models (LLMs), which can adeptly replicate these features. This challenge underscores the need for enhanced methodologies that not only differentiate between human and AI-generated content but also explore how LLMs emulate specific authorial styles (Boenninghoff et al., 2019a; Danilevsky et al., 2020).

### v2

@@ 5. Human-LLM Co-authored Text Attribution
Besides creating text from scratch, LLMs are often used for extending sequences from human prompts. These perturbations, have diminished the effectiveness of existing text detection methods (Bhat and Parthasarathy, 2020). Identifying text that combines input from both human authors and LLMs presents unique challenges. Hybrid texts may originate as human-written content, with LLMs employed for conditional generation, making it difficult to clearly distinguish between the stylistic features of human and machine contributions. There are fewer studies on this task due to its difficulty, and existing research often makes simplifications.
@@ 5.2. Methodologies
To effectively analyze and classify human-LLM co-authored texts versus those solely authored by humans or LLMs, feature-based methods from LLM-generated text detection, such as Log-likelihood (Solaiman et al., 2019), GLTR (Gehrmann et al., 2019), and log-rank (Mitchell et al., 2023), are adapted to this task. Additionally, neural network-based models like BERT (Ippolito et al., 2020), Radar (Hu et al., 2023), and GPT-sentinel (Chen et al., 2023a) can also be applied. The complexity of this task increases as users may employ multiple LLMs to compose different sections of an article, further blurring the lines between human and machine-generated content. Consequently, techniques used in earlier LLM detection need to evolve continuously. This ongoing evolution in detection strategies mirrors the increasing sophistication of LLM outputs and the collaborative nature of modern text creation.
@@ 5.3. Open Challenges
Human-LLM co-authored texts pose a more intricate challenge due to the blending of stylistic and linguistic elements from both human authors and LLMs. These texts may begin as human drafts later extended or revised by LLMs, or vice versa, resulting in an integration of styles that obscures distinct authorial markers (Liu et al., 2023a). Advanced techniques are required to dissect these integrations, identifying where and how LLM contributions intersect with human input (Wang et al., 2024). This involves detecting subtle shifts in style and contextual cues that indicate the extent of LLM involvement, allowing for accurate segmentation and attribution of authorship within hybrid documents.
@@ 7.2. Generalization
This subsection examines the applicability of current methodologies across varying LLMs, domains, genres, and languages. Domains refer to broad areas of knowledge or topics, while genres refer to specific styles or forms of writing within any domain. Domain generalization poses significant challenges due to variability in vocabularies, syntax, and styles across different subjects, complicating accurate authorship attribution. Attribution performance tends to drop when known and query texts differ in topic or genre (Altakrori et al., 2021). Models like BERT and RoBERTa have shown limitations in cross-domain tasks (Barlas and Stamatatos, 2020; Huertas-Tato et al., 2022), and adapting models to new domains remains difficult due to factors like dataset size variability and writing styles. Traditional methods focused on identifying less topic-dependent features such as function words and part-of-speech n-grams (Madigan et al., 2005b; Menon and Choi, 2011), while recent approaches highlight the importance of training more powerful transformer-based models (Rivera-Soto et al., 2021) and techniques such as adversarial training (Li et al., 2017; Ben-David et al., 2010; Ganin et al., 2016).
@@
Improving generalization can potentially be achieved through several strategies. First, leveraging transfer learning by pre-training on large, diverse datasets and fine-tuning on specific domains enhances adaptability and performance (Barlas and Stamatatos, 2020; Rodriguez et al., 2022). Second, developing domain and genre-invariant features would improve robustness by focusing on core stylistic elements (Argamon et al., 2003). Third, employing hybrid models and ensemble methods that integrate domain-specific knowledge can optimize prediction accuracy by drawing on the strengths of individual models (Bacciu et al., 2019). Additionally, incorporating contextual factors like the writing environment or intended audience, alongside data augmentation techniques, can bolster generalization. Finally, improving attribution in multilingual contexts enables models to operate effectively across various languages (Chen et al., 2022; Shamardina et al., 2022). As detectors could be biased against non-native English writers (Liang et al., 2023), enhancing multilingual generalization is crucial for fairness. Collectively, these approaches foster robust and adaptable models equipped to handle diverse styles and contexts.
@@ 7.3. Explainability
Improving explainability is crucial for ensuring transparency and trust in authorship attribution models, especially as they become increasingly integrated into fields such as forensic investigation, academia, and journalism. Developing explainable AI techniques specifically for authorship attribution can lead to methodologies where the reasoning behind attributions is clear and understandable. This transparency is particularly important when authorship attribution is used as evidence in legal proceedings (Chaski, 2005; Rocha et al., 2016). Traditional attribution methods, which rely on stylistic and linguistic features to identify an author, often struggle to distinguish between human-authored texts and those generated by LLMs, which can adeptly replicate these features. This challenge underscores the need for enhanced methodologies that not only differentiate between human and AI-generated content but also explore how LLMs emulate specific authorial styles (Boenninghoff et al., 2019a; Danilevsky et al., 2020).
@@ 9. Conclusion
Acknowledgments
This material is based upon work supported by the U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-04, NSF awards (SaTC-2241068, IIS-2339198, and POSE-2346158), a Cisco Research Award, and a Microsoft Accelerate Foundation Models Research Award. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security and the National Science Foundation.